# Efficient wireless charging of a quantum battery

### v1

@@ IV.1 The case of a single charging unit
In the bottom two panels of Fig. 4, we show the $R$ dependence of the ergotropy $\bar{\mathcal{E}}$ charged on the QB at $t=\bar{t}$. One can see that for both the scenarios I and II, $\bar{\mathcal{E}}$ increases monotonically with the increase of $R$. When $R\rightarrow\infty$, $\bar{\mathcal{E}}$ approaches its asymptotic value, which, for the scenario I, equals $c_{1}\omega_{0}$, and for the scenario II, equals $\omega_{0}$. This indicates that in theory, the energy in the charger is almost fully transferred to extractable work in the QB for such a limiting case. In fact, if the charger is in the excited state and QB is in the ground state, even for the $R\simeq 10$ case which is experimentally accessible apl , the QB can be charged up to ¯$\bar{\mathcal{E}}\simeq 0.72\omega_{0}$.
@@
When $R$ is finite, the ergotropy $\bar{\mathcal{E}}$ charged on a QB depends on $c_{1}$ and $e_{1}$. To elucidate this in detail, we show in Fig. 5 the $c_{1}$ dependence of $\bar{\mathcal{E}}$ with the QB being in its ground state initially and in Fig. 6 the $e_{1}$ dependence of $\bar{\mathcal{E}}$ with the charger being in the excited state initially. For the scenario I, when $c_{1}=0$, ¯$\bar{\mathcal{E}}=0$ by definition. When $c_{1}$ increases from 0 to 1, as depicted in Fig. 5(a), $\bar{\mathcal{E}}$ first increases to a small peak and then decreases to zero (e.g., $R=0.1$ and $1$) or to a small but finite value (e.g., $R=2$) in the weak coupling regime. In the strong coupling regime, however, $\bar{\mathcal{E}}$ increases monotonically from 0 to its maximum; in particular, for large enough $R$ (e.g., $R=100$), it increases nearly linearly with the increase of $c_{1}$. Moreover, by comparing the three panels in Fig. 5, one can also see that the coherent contribution to ergotropy is dominant when $c_{1}$ is relatively small, while this is not the case for large enough $c_{1}$, and this phenomenon is consistent with that shown in Fig. 2. In particular, if the charger is in the excited state (i.e., $c_{1}=1$) initially, ¯¯$\bar{\mathcal{E}}\equiv\bar{\mathcal{E}}_{i}$. This is because for this case, the battery state $\rho_{\mathrm{ba}}(\bar{t})$ at time $\bar{t}$ is diagonal, thereby there is no coherent contribution to the ergotropy shihl2 .
@@
For the scenario II, the QB is not fully discharged at the initial time, that is, there is residual ergotropy in it. For this case, when $R$ is very small (i.e., the bad cavity limit), the QB cannot be charged, so we plot in Fig. 6 the case of $R\geqslant 2$, from which one can see that even for $R=2$, the charged ergotropy $\bar{\mathcal{E}}$ is still very small, and it exceeds the initial ergotropy $e_{1}\omega_{0}$ of the QB only in a very narrow region of $e_{1}$ (i.e., $e_{1}\lesssim 0.0382$). By increasing the coupling strength of the qubits and reservoir, $\bar{\mathcal{E}}$ can be noticeably enhanced. From Fig. 6(a) one can also see that the region of $e_{1}$ in which $\bar{\mathcal{E}}$ is larger than the initial ergotropy $e_{1}\omega_{0}$ expands with the increasing $R$. For example, for $R=100$, the associated region can be obtained approximately as $e_{1}\lesssim 0.9239$. As for the ratio of the incoherent and coherent components in $\bar{\mathcal{E}}$, from the bottom two panels of Fig. 6, one can see that apart from the small $R$ case (e.g., $R=2$), the incoherent contribution is always dominant, and this is different from that for the scenario I. We would also like to mention that $\bar{\mathcal{E}}$ shown in Fig. 6(a) is not always a monotonic decreasing function of $e_{1}$ (or equivalently, the initial ergotropy in the QB), e.g., when $R=100$ it takes a minimum at $e_{1}\simeq 0.5636$. Nonetheless, $\bar{\mathcal{E}}$ always takes its maximum at $e_{1}=0$, that is, the QB is in its ground state and empty initially. We have also examined the mixed battery state $\rho_{\mathrm{ba}}=\mathrm{diag}\{e_{1},1-e_{1}\}$, which is empty for $e_{1}\leqslant 0.5$, although its energy is $e_{1}\omega_{0}$. It is found that the charging performance is less efficient than that prepared in the ground state. In this sense, the optimal charging of a QB is achieved starting from its ground state, which is in agreement with the findings of Ref. initial .
@@
Up to now, we have elucidated how the enlarged $R$ helps to accelerate the charging rate and to pump more ergotropy (i.e., the maximal amount of extractable work) in the QB. Here, it is natural to ask another intriguing question: is the mean energy or ergotropy of the charger determines the ergotropy charged on a QB? As they always coexist for the free Hamiltonian $H_{\mathrm{ch}}$ [see Eq. (7)] and the initial state $|\psi\rangle_{c_{1}}$ [see Eq. (11)], we turn to consider a general one-qubit charger state $\rho_{\mathrm{ch}}$ and we first figure out the condition under which there is nonzero energy and zero ergotropy in it. After some algebra, one can obtain that this is achieved for $\rho_{\mathrm{ch}}=\mathrm{diag}\{c_{1},1-c_{1}\}$ with $c_{1}\leqslant 0.5$. In this case, the mean energy $c_{1}\omega_{0}$ in the charger equals to that for the initial state $|\psi\rangle_{c_{1}}$, and in the good cavity limit, most of the energy can be pumped in the QB. However, it is found that for any initial state (including the mixed one) of the QB, there is no ergotropy being charged on it. In this sense, it seems that it is the charger ergotropy instead of its energy determines the ergotropy charged on a QB.
@@
To gain more insight into the wireless charging process of the QB, we can further consider dynamics of the mean energy $E_{\mathrm{ch}}$ and ergotropy $\mathcal{E}_{\mathrm{ch}}$ in the charger. Physically, as there is no direct interaction between the charger and the QB, the energy (ergotropy) in the charger will first being pumped in the reservoir and then charged into the QB via the reservoir-mediated indirect interaction. During such a dynamical process, partial of the energy (ergotropy) will inevitably be lost. Here, we focus on the case of not very small $R$ and calculate the ergotropy in the charger with the same parameters as in Fig. 2 (for the conciseness of this paper, we do not show them here). For the scenario I, it is found that the ergotropy of the charger reaches its first dynamical minimum at $t=\bar{t}$, and such a minimum decreases with the increase of $R$, for example, for $R=20$ it is of approximately 0. For the scenario II, however, the ergotropy $\mathcal{E}$ in the QB does not always increase synchronously with the decrease of the ergotropy $\mathcal{E}_{\mathrm{ch}}$ in the charger. Specifically, the ergotropy for both the charger and the QB decays to zero after a short period of evolution time and then turns to be increased to the first dynamical maxima. However, the former decays to zero faster than the latter. This shows that there is a delayed effect for the ergotropy pumped in the reservoir to be charged into the QB.
@@ IV.2 The case of two charging units
First, we show in Fig. 7 time evolution of $\mathcal{E}$, $\mathcal{E}_{i}$, and $\mathcal{E}_{c}$ of the QB in the good cavity limit $R=20$, where we have chosen $c_{1}=c_{2}$ and the values of the system parameters are the same as those in Fig. 2. From this figure one can note that for both the scenarios I and II, $\mathcal{E}$ shows a structurally similar behavior to that with a single charging unit. Moreover, although in this case the initial energy and ergotropy in the charger equal twice those for a single charging unit, the dynamical maximum $\bar{\mathcal{E}}$ is slightly decreased. This implies that the amounts of initial energy and ergotropy in the charger are not the only ingredients determining the ergotropy charged on a QB.
@@
The incoherent and coherent components of ergotropy also show similar behaviors to those obtained for the case of a single charging unit shown in Fig. 2. Specifically, for the scenario I, the incoherent (coherent) contribution to ergotropy is dominant for the charger with a large (small) $c_{1,2}$, while for the scenario II, the incoherent contribution to ergotropy is always dominant. Looking at Fig. 7, one can also see that both $\mathcal{E}$ and $\mathcal{E}_{i}$ reach their first dynamical maxima at the same interaction time $\bar{t}$, but this is not the case for $\mathcal{E}_{c}$ in general.
@@
In the bad cavity limit $R=0.1$, as exemplified in the bottom two panels of Fig. 3, the ergotropy $\mathcal{E}$ also shows qualitatively the same behavior to that for the case of a single charging unit. But for the scenario I, apart from the special case of $c_{1,2}=1$ for which $\mathcal{E}$ always remains zero, the dynamical maximum of $\mathcal{E}$ in this case is noticeably enhanced, and this phenomenon is in sharp contrast to that in the good cavity limit.
@@
Figure 8: The $R$ dependence of the charging time $\lambda\bar{t}$ and ergotropy $\bar{\mathcal{E}}$ charged by two charging units at time $\bar{t}$ The left (right) two panels are plotted with the charger and QB being in the initial states $|\psi\rangle_{c_{1}c_{2}}$ and $|0\rangle$ ($|11\rangle$ and $|\varphi\rangle_{e_{1}}$), respectively. For small $R$, $\lambda\bar{t}$ is very large and the top two panels are cut to better visual its behavior in the large $R$ region, and the first five points in (c) and (d) are plotted with $R=0.1$, 1, 2, 3 and 4, respectively.
@@
In Fig. 8, we show the $R$ dependence of the charging time $\bar{t}$ and the associated $\bar{\mathcal{E}}$ with different system parameters. First, one can see that its behavior is very similar to that for a single charging unit (cf. Fig. 4). Specifically, it is also insensitive to the variation of $c_{1,2}$ (scenario I) and $e_{1}$ (scenario II), especially in the strong coupling regime. By comparing Figs. 4 and 8, one can also find that $\lambda\bar{t}$ is slightly shortened in this case. Moreover, while $\bar{\mathcal{E}}$ also approaches an asymptotic value when $R\rightarrow\infty$, in contrast to the case of a single charging unit, such an asymptotic value is smaller than the initial energy $2c_{1}\omega_{0}$ (equals to the initial ergotropy) in the two charging units. In particular, for the scenario II with $e_{1}\gtrsim 0.5911$ [see, e.g., the triangles in Fig. 8(d)], $\bar{\mathcal{E}}$ even cannot exceed its initial value $e_{1}\omega_{0}$, that is, the QB in this case cannot be charged anymore by the two charging units. This indicates that in the good cavity limit, the two charging units cannot outperform the single charging one if there is considerable amount of residual ergotropy in the QB initially. Of course, the case may be different for the finite $R$, and we will discuss this issue in detail after introducing the three charging units.
@@
Figure 9: Comparison of $\bar{\mathcal{E}}$ charged on a QB in the ground state initially, where the charger is in the initial state $|\psi\rangle_{c_{1}c_{1}}$ (circles) or $|\Psi^{+}\rangle$ (squares). The left dashed lines correspond to $c_{1}=0.5$ after which the charger energy $E_{\mathrm{ch}}(|\Psi^{+}\rangle)<E_{\mathrm{ch}}(|\psi\rangle_{c_{1}c_{1}})$, while the right dashed lines at $c_{1}\simeq 0.6045$ (a) and $0.6575$ (b), respectively, show the critical points at which $\bar{\mathcal{E}}$ for $|\Psi^{+}\rangle$ equals to that for $|\psi\rangle_{c_{1}c_{1}}$. The star in (b) denotes the case for which $\bar{\mathcal{E}}$ corresponds to the second instead of the first dynamical maximum of $\mathcal{E}$.
@@
As we studied in the above only the initial product charger states, it is natural to ask whether the charging performance of a QB can be improved by using the initially quantum correlated charger states. To answer this question, we consider the following Bell-like charger states:
@@
and in Fig. 9 we give a plot of $\bar{\mathcal{E}}$ charged on a QB in its ground state by using the charger in the initial states $|\psi\rangle_{c_{1}c_{1}}$ and $|\Psi^{+}\rangle$, respectively. The initial charger energy is $E_{\mathrm{ch}}(|\psi\rangle_{c_{1}c_{1}})=2c_{1}\omega_{0}$ and $E_{\mathrm{ch}}(|\Psi^{+}\rangle)=\omega_{0}$ ($\forall c_{1}$), respectively, and so is the initial charger ergotropy. As shown in Fig. 9, in the parameter region between the two dashed vertical lines, $\bar{\mathcal{E}}$ for the initially correlated charger state is larger than that for the initially product charger state, though for the former case there is less energy in the charger. This indicates that, in certain situations, the correlated charger states are more efficient than that of the uncorrelated ones. But this is not a general conclusion; for example, a further calculation shows that the charger states $|\Psi^{-}\rangle$ and $|\Phi^{\pm}\rangle$ are less efficient than $|\psi\rangle_{c_{1}c_{1}}$. Moreover, as highlighted by the star in Fig. 9(b), the maximum ergotropy $\bar{\mathcal{E}}$ charged on a QB might correspond to the second instead of the first dynamical maximum of $\mathcal{E}$.
@@
We have also examined time evolution of the ergotropy $\mathcal{E}_{\mathrm{ch}}$ in the two charging units, and for the conciseness of this paper, we also do not display them here. Note that $\mathcal{E}_{\mathrm{ch}}$ represents the ergotropy of the charger with respect to the total state $\rho_{\mathrm{ch}}$ of the two charging units, which is larger than the sum of the ergotropies with respect to the reduced states of the two charger qubits EPL . Similar to that for a single charging unit, the interaction time at which $\mathcal{E}_{\mathrm{ch}}$ reaches its first dynamical minimum is also not synchronously with the charging time $\bar{t}$ at which the ergotropy charged on the QB reaches its dynamical maximum. Specifically, for the scenario I with not very small $R$, $\mathcal{E}_{\mathrm{ch}}$ reaches its first dynamical minimum after $t=\bar{t}$ if $c_{1,2}$ is very large (e.g., $c_{1,2}\gtrsim 0.9756$ when $R=20$), while for the other $c_{1,2}$, it reaches its first dynamical minimum before $t=\bar{t}$. For the scenario II with not very small $R$, however, $\mathcal{E}_{\mathrm{ch}}$ reaches its first dynamical minimum after $t=\bar{t}$ for a very small $e_{1}$ (e.g., $e_{1}\lesssim 0.0476$ when $R=20$), and otherwise, it reaches its first dynamical minimum before $t=\bar{t}$. This indicates that the energy pumped in the reservoir previously can still be transferred to the QB in the time interval during which there is no additional energy to be injected into the reservoir.

### v2

@@ IV.1 The case of a single charging unit
In the bottom two panels of Fig. 4, we show the $R$ dependence of the ergotropy $\bar{\mathcal{E}}$ charged on the QB at $t=\bar{t}$. One can see that for both scenarios I and II, $\bar{\mathcal{E}}$ increases monotonically with the increase of $R$. When $R\rightarrow\infty$, $\bar{\mathcal{E}}$ approaches its asymptotic value, which, for scenario I, equals $c_{1}\omega_{0}$, and for scenario II, equals $\omega_{0}$. This indicates that in theory, the energy in the charger is almost fully transferred to extractable work in the QB for such a limiting case. In fact, if the charger is in the excited state and QB is in the ground state, even for the $R\simeq 10$ case which is experimentally accessible apl , the QB can be charged up to ¯$\bar{\mathcal{E}}\simeq 0.72\omega_{0}$.
@@
When $R$ is finite, the ergotropy $\bar{\mathcal{E}}$ charged on a QB depends on $c_{1}$ and $e_{1}$. To elucidate this in detail, we show in Fig. 5 the $c_{1}$ dependence of $\bar{\mathcal{E}}$ with the QB being in its ground state initially, and in Fig. 6 the $e_{1}$ dependence of $\bar{\mathcal{E}}$ with the charger being in the excited state initially. For scenario I, when $c_{1}=0$, ¯$\bar{\mathcal{E}}=0$ by definition. When $c_{1}$ increases from 0 to 1, as depicted in Fig. 5(a), $\bar{\mathcal{E}}$ first increases to a small peak and then decreases to zero (e.g., $R=0.1$ and $1$) or to a small but finite value (e.g., $R=2$) in the weak coupling regime. In the strong coupling regime, however, $\bar{\mathcal{E}}$ increases monotonically from 0 to its maximum; in particular, for large enough $R$ (e.g., $R=100$), it increases nearly linearly with the increase of $c_{1}$. Moreover, by comparing the three panels in Fig. 5, one can also see that the coherent contribution to ergotropy is dominant when $c_{1}$ is relatively small, while this is not the case for large enough $c_{1}$, and this phenomenon is consistent with that shown in Fig. 2. In particular, if the charger is in the excited state (i.e., $c_{1}=1$) initially, ¯¯$\bar{\mathcal{E}}\equiv\bar{\mathcal{E}}_{i}$. This is because for this case, the battery state $\rho_{\mathrm{ba}}(\bar{t})$ at time $\bar{t}$ is diagonal, thereby there is no coherent contribution to the ergotropy shihl2 .
@@
For scenario II, the QB is not fully discharged at the initial time, that is, there is residual ergotropy in it. For this case, when $R$ is very small (i.e., the bad cavity limit), the QB cannot be charged, so we plot in Fig. 6 the case of $R\geqslant 2$, from which one can see that even for $R=2$, the charged ergotropy $\bar{\mathcal{E}}$ is still very small, and it exceeds the initial ergotropy $e_{1}\omega_{0}$ of the QB only in a very narrow region of $e_{1}$ (i.e., $e_{1}\lesssim 0.0382$). By increasing the coupling strength of the qubits and reservoir, $\bar{\mathcal{E}}$ can be noticeably enhanced. From Fig. 6(a) one can also see that the region of $e_{1}$ in which $\bar{\mathcal{E}}$ is larger than the initial ergotropy $e_{1}\omega_{0}$ expands with increasing $R$. For example, for $R=100$, the associated region can be obtained approximately as $e_{1}\lesssim 0.9239$. As for the ratio of the incoherent and coherent components in $\bar{\mathcal{E}}$, from the bottom two panels of Fig. 6, one can see that apart from the small $R$ case (e.g., $R=2$), the incoherent contribution is always dominant, and this is different from that for scenario I. We would also like to mention that $\bar{\mathcal{E}}$ shown in Fig. 6(a) is not always a monotonic decreasing function of $e_{1}$ (or equivalently, the initial ergotropy in the QB), e.g., when $R=100$ it takes a minimum at $e_{1}\simeq 0.5636$. Nonetheless, $\bar{\mathcal{E}}$ always takes its maximum at $e_{1}=0$, that is, the QB is in its ground state and empty initially. We have also examined the mixed battery state $\rho_{\mathrm{ba}}=\mathrm{diag}\{e_{1},1-e_{1}\}$, which is empty for $e_{1}\leqslant 0.5$, although its energy is $e_{1}\omega_{0}$. It is found that the charging performance is less efficient than that prepared in the ground state. In this sense, the optimal charging of a QB is achieved starting from its ground state, which is in agreement with the findings of Ref. initial .
@@
Up to now, we have elucidated how the enlarged $R$ helps to accelerate the charging rate and to pump more ergotropy (i.e., the maximal amount of extractable work) in the QB. Here, it is natural to ask another intriguing question: does the mean energy or ergotropy of the charger determine the ergotropy charged on a QB? As they always coexist for the free Hamiltonian $H_{\mathrm{ch}}$ [see Eq. (7)] and the initial state $|\psi\rangle_{c_{1}}$ [see Eq. (11)], we turn to consider a general one-qubit charger state $\rho_{\mathrm{ch}}$ and we first figure out the condition under which there is nonzero energy and zero ergotropy in it. After some algebra, one can obtain that this is achieved for $\rho_{\mathrm{ch}}=\mathrm{diag}\{c_{1},1-c_{1}\}$ with $c_{1}\leqslant 0.5$. In this case, the mean energy $c_{1}\omega_{0}$ in the charger equals that for the initial state $|\psi\rangle_{c_{1}}$, and in the good cavity limit, most of the energy can be pumped in the QB. However, it is found that for any initial state (including the mixed one) of the QB, there is no ergotropy being charged on it. In this sense, it seems that it is the charger ergotropy instead of its energy that determines the ergotropy charged on a QB.
@@
To gain more insight into the wireless charging process of the QB, we can further consider the dynamics of the mean energy $E_{\mathrm{ch}}$ and ergotropy $\mathcal{E}_{\mathrm{ch}}$ in the charger. Physically, as there is no direct interaction between the charger and the QB, the energy (ergotropy) in the charger will first be pumped in the reservoir and then charged into the QB via the reservoir-mediated indirect interaction. During such a dynamical process, part of the energy (ergotropy) will inevitably be lost. Here, we focus on the case of not very small $R$ and calculate the ergotropy in the charger with the same parameters as in Fig. 2 (for the conciseness of this paper, we do not show them here). For scenario I, it is found that the ergotropy of the charger reaches its first dynamical minimum at $t=\bar{t}$, and such a minimum decreases with the increase of $R$, for example, for $R=20$ it is of approximately 0. For scenario II, however, the ergotropy $\mathcal{E}$ in the QB does not always increase synchronously with the decrease of the ergotropy $\mathcal{E}_{\mathrm{ch}}$ in the charger. Specifically, the ergotropy for both the charger and the QB decays to zero after a short period of evolution time and then turns to be increased to the first dynamical maxima. However, the former decays to zero faster than the latter. This shows that there is a delayed effect for the ergotropy pumped in the reservoir to be charged into the QB.
@@ IV.2 The case of two charging units
First, we show in Fig. 7 the time evolution of $\mathcal{E}$, $\mathcal{E}_{i}$, and $\mathcal{E}_{c}$ of the QB in the good cavity limit $R=20$, where we have chosen $c_{1}=c_{2}$ and the values of the system parameters are the same as those in Fig. 2. From this figure one can note that for both scenarios I and II, $\mathcal{E}$ shows a structurally similar behavior to that with a single charging unit. Moreover, although in this case the initial energy and ergotropy in the charger equal twice those for a single charging unit, the dynamical maximum $\bar{\mathcal{E}}$ is slightly decreased. This implies that the amounts of initial energy and ergotropy in the charger are not the only ingredients determining the ergotropy charged on a QB.
@@
The incoherent and coherent components of ergotropy also show similar behaviors to those obtained for the case of a single charging unit shown in Fig. 2. Specifically, except the case of $c_{1}=1$ and $e_{1}=0$ for which $\mathcal{E}_{c}\equiv 0$, for scenario I (II) with $c_{1}\gtrsim 0.5078$ ($e_{1}\neq 0$ and 1), the two components of ergotropy dominate its dynamics alternately, whereas for scenario I with $c_{1}\lesssim 0.5078$, the coherent contribution is always dominant. From Fig. 7 one can also see that both $\mathcal{E}$ and $\mathcal{E}_{i}$ reach their first dynamical maxima at the same interaction time $\bar{t}$, but this is not the case for $\mathcal{E}_{c}$ in general.
@@
In the bad cavity limit $R=0.1$, as exemplified in the bottom two panels of Fig. 3, the ergotropy $\mathcal{E}$ also shows qualitatively the same behavior as that for the case of a single charging unit. But for scenario I, apart from the special case of $c_{1,2}=1$ for which $\mathcal{E}$ always remains zero, the dynamical maximum of $\mathcal{E}$ in this case is noticeably enhanced, and this phenomenon is in sharp contrast to that in the good cavity limit.
@@
Figure 8: The $R$ dependence of the charging time $\lambda\bar{t}$ and ergotropy $\bar{\mathcal{E}}$ charged by two charging units at time $\bar{t}$ The left (right) two panels are plotted with the charger and QB being in the initial states $|\psi\rangle_{c_{1}c_{2}}$ and $|0\rangle$ ($|11\rangle$ and $|\varphi\rangle_{e_{1}}$), respectively. For small $R$, $\lambda\bar{t}$ is very large and the top two panels are cut to better see its behavior in the large $R$ region, and the first five points in (c) and (d) are plotted with $R=0.1$, 1, 2, 3, and 4, respectively.
@@
In Fig. 8, we show the $R$ dependence of the charging time $\bar{t}$ and the associated $\bar{\mathcal{E}}$ with different system parameters. First, one can see that its behavior is very similar to that for a single charging unit (cf. Fig. 4). Specifically, it is also insensitive to the variation of $c_{1,2}$ (scenario I) and $e_{1}$ (scenario II), especially in the strong coupling regime. This is also due to the fact that the indirect coupling of the charger and QB is induced by the cavity mode through the exchange of cavity photons, thus the decay rate of the pseudomode is determined by the coupling strength $\Omega$ (note that $R=\sqrt{2}\Omega/\lambda$) common1 ; common2 . By comparing Figs. 4 and 8, one can also find that $\bar{t}$ is slightly shortened in this case. Moreover, while $\bar{\mathcal{E}}$ also approaches an asymptotic value when $R\rightarrow\infty$, in contrast to the case of a single charging unit, such an asymptotic value is smaller than the initial energy $2c_{1}\omega_{0}$ (equals the initial ergotropy) in the two charging units. In particular, for scenario II with $e_{1}\gtrsim 0.5911$ [see, e.g., the triangles in Fig. 8(d)], $\bar{\mathcal{E}}$ even cannot exceed its initial value $e_{1}\omega_{0}$, that is, the QB in this case cannot be charged anymore by the two charging units. This indicates that in the good cavity limit, the two charging units cannot outperform the single charging one if there is a considerable amount of residual ergotropy in the QB initially. Of course, the case may be different for the finite $R$, and we will discuss this issue in detail after introducing the three charging units.
@@
Figure 9: Comparison of $\bar{\mathcal{E}}$ charged on a QB in the ground state initially, where the charger is in the initial state $|\psi\rangle_{c_{1}c_{1}}$ (circles) or $|\Psi^{+}\rangle$ (squares). The left dashed lines correspond to $c_{1}=0.5$, after which the charger energy $E_{\mathrm{ch}}(|\Psi^{+}\rangle)<E_{\mathrm{ch}}(|\psi\rangle_{c_{1}c_{1}})$, while the right dashed lines at $c_{1}\simeq 0.6045$ (a) and $0.6575$ (b), respectively, show the critical points at which $\bar{\mathcal{E}}$ for $|\Psi^{+}\rangle$ equals that for $|\psi\rangle_{c_{1}c_{1}}$. The star in (b) denotes the case for which $\bar{\mathcal{E}}$ corresponds to the second instead of the first dynamical maximum of $\mathcal{E}$.
@@
As we studied in the above only the initial product charger states, it is natural to ask whether the charging performance of a QB can be improved by using the initially correlated charger state. To answer this question, we consider the following Bell-like charger states:
@@
which have the same entanglement. In Fig. 9, we give a plot of $\bar{\mathcal{E}}$ charged on a QB in its ground state by using the charger in the initial states $|\psi\rangle_{c_{1}c_{1}}$ and $|\Psi^{+}\rangle$, respectively. The initial charger energy is $E_{\mathrm{ch}}(|\psi\rangle_{c_{1}c_{1}})=2c_{1}\omega_{0}$ and $E_{\mathrm{ch}}(|\Psi^{+}\rangle)=\omega_{0}$ ($\forall c_{1}$), respectively, and so is the initial charger ergotropy. As a result, $E_{\mathrm{ch}}(|\Psi^{+}\rangle)<E_{\mathrm{ch}}(|\psi\rangle_{c_{1}c_{1}})$ when $c_{1}>0.5$. As is shown in Fig. 9, the charged ergotropy with the initial charger state $|\Psi^{+}\rangle$ decreases with the increasing $|c_{1}-0.5|$, which is proportional to the initial entanglement in it. Moreover, the curves of ergotropy $\bar{\mathcal{E}}$ for these two initial charger states cross each other at a critical point $c_{1,r}$. In the region between 0 and $c_{1,r}$, $\bar{\mathcal{E}}$ for the initial charger state $|\Psi^{+}\rangle$ is larger than that for the initial charger state $|\psi\rangle_{c_{1}c_{1}}$; in particular, it is worth emphasizing that when $c_{1}\in(0.5,c_{1,r})$, the mean energy in $|\Psi^{+}\rangle$ is less than that in $|\psi\rangle_{c_{1}c_{1}}$. This indicates that, under certain situations, the correlated charger states are more efficient than that of the uncorrelated ones. But, this is not a universal conclusion. For example, ¯$\bar{\mathcal{E}}\equiv 0$ for the initial charger state $|\Psi^{-}\rangle$. This is understandable as it is robust against the decoherence effect; in particular, it reduces to the decoherence-free state when $c_{1}=0.5$ common1 ; common2 , for which there is no energy being transferred to the QB. For the initial charger states $|\Phi^{\pm}\rangle$, a further calculation reveals that they are also less efficient than $|\psi\rangle_{c_{1}c_{1}}$, although they have the same entanglement with $|\Psi^{+}\rangle$ and there is a considerable amount of energy ($E_{\mathrm{ba}}\simeq 0.8799c_{1}\omega_{0}$) being transferred to the QB. Thus, the advantage of the entangled states depends on the structure of the entanglement, the spectral properties of the reservoir, and the coupling regime. Moreover, as highlighted by the star in Fig. 9(b), the maximum ergotropy $\bar{\mathcal{E}}$ charged on the QB may correspond to the second instead of the first dynamical maximum of $\mathcal{E}$.
@@
We have also examined time evolution of the ergotropy $\mathcal{E}_{\mathrm{ch}}$ in the two charging units, and for the conciseness of this paper, we also do not display them here. Note that $\mathcal{E}_{\mathrm{ch}}$ represents the ergotropy of the charger with respect to the total state $\rho_{\mathrm{ch}}$ of the two charging units, which is larger than the sum of the ergotropies with respect to the reduced states of the two charger qubits EPL . Similar to that for a single charging unit, the interaction time at which $\mathcal{E}_{\mathrm{ch}}$ reaches its first dynamical minimum is also not synchronous with the charging time $\bar{t}$ at which the ergotropy charged on the QB reaches its dynamical maximum. Specifically, for scenario I with not very small $R$, $\mathcal{E}_{\mathrm{ch}}$ reaches its first dynamical minimum after $t=\bar{t}$ if $c_{1,2}$ is very large (e.g., $c_{1,2}\gtrsim 0.9756$ when $R=20$), while for the other $c_{1,2}$, it reaches its first dynamical minimum before $t=\bar{t}$. For scenario II with not very small $R$, however, $\mathcal{E}_{\mathrm{ch}}$ reaches its first dynamical minimum after $t=\bar{t}$ for a very small $e_{1}$ (e.g., $e_{1}\lesssim 0.0476$ when $R=20$), and otherwise, it reaches its first dynamical minimum before $t=\bar{t}$. This indicates that the energy pumped in the reservoir previously can still be transferred to the QB in the time interval during which there is no additional energy to be injected into the reservoir.